OpenStudy (anonymous):

find the quadratic function y=ax^2+bx+c whose graph passes through the given points. (-1,14) (1,4) (3,2)

OpenStudy (anonymous):

dear u should put your data in your function i mean : 14=a(-1)^2+(-1x+c 4=a1^2+b1+c 2=a3^2+b3+c now you have 3 equation and 3 unknown parameter so u can solve it .